OpenStudy (anonymous):

I really need help ive done this but I don't remember how to do this factor 2x^2+8x-90

OpenStudy (anonymous):

(2x+18)(x-5)

OpenStudy (anonymous):

First, I notice that all of the terms are divisible by 2 2x^2+8x-90 = 2(x^2+4x-45) Then we need to think, what times what makes -45, but adds together to make 4. Knowing this we need one value to be negative and one positive. As well, we need the positive value to be larger than the negative. I think, well 9*5 makes 45, so if we do 9 * -5, that makes 45, and also 9+ (-5) = 4 so: 2(x^2+4x-45)= 2(x+9)(x-5)

OpenStudy (anonymous):

Alternatively, after we factor out the 2, we can plug it into the quadratic equation. \[\frac{ -b+\sqrt{b^2-4ac} }{ 2a }\] and \[\frac{ -b-\sqrt{b^2-4ac} }{ 2a }\] for when \[ax^2+bx+c=0\]

OpenStudy (anonymous):

thank you, it clicked back into my memorie thank you so much that was so much help